OpenStudy (anonymous):

Thirty-six percent of what number is 180? (Points : 4) 500 515 540 550

OpenStudy (anonymous):

Well, to start, how do you write 36 percent as a decimal?

OpenStudy (anonymous):

you could do cross multiplication so... 36/100 = x/180 Thats how you would start out ... That would be equal to 36x=100(180) Which is simplified to 36x=18000 You would then divide by 36 on BOTH sides of the equation. It would equal 500! So 36% of 500 = 180 :D